OpenStudy (anonymous):

Find the area of a rectangle with base of 12 inches and height of 8 inches

OpenStudy (ja1):

A = lw

OpenStudy (ja1):

so 12 * 8 =

OpenStudy (ja1):

do you understand?

OpenStudy (anonymous):

A= l * w = 12 * 8 = 96

OpenStudy (anonymous):

Rectangles involve using this formula to find the area: Area = Length x Width